OpenStudy (anonymous):

Simplify (1 − sin x)(1 + sin x).

OpenStudy (luigi0210):

FOIL it out.

OpenStudy (luigi0210):

If you do it right, it'll give you a trig identity

OpenStudy (anonymous):

Use (a-b)(a+b) = a^2-b^2 and sin^2 + cos^2 = 1 ?

OpenStudy (anonymous):

(1- sin x)^2

OpenStudy (anonymous):

that is what I got...

OpenStudy (anonymous):

(a-b)(a+b) does not equal (a-b)^2

OpenStudy (anonymous):

1^2 - sin x ^2 which is 1 - sin x ^2 which is what I got when I FOILED it...

OpenStudy (johnweldon1993):

Well correct but remember it will be \[\large 1^2 - sin^2(x)\] so what is \[\large 1 - \sin^2(x) = \space ?\]

OpenStudy (johnweldon1993):

Hint would be to remember the identity \[\large \cos^2(x) + \sin^2(x) = 1\]

OpenStudy (anonymous):

Yes, I actually have that identity written down. But I am not sure how they tie in together in this problem.

OpenStudy (johnweldon1993):

Well...if \[\large cos^2(x) + sin^2(x) = 1\] what would happen if you subtracted \(\large sin^2(x)\) from both sides? \[\large 1 - sin^2(x) = cos^2(x)\]

OpenStudy (anonymous):

cos^2 (x) = 1 - sin^2(x)

OpenStudy (anonymous):

And you already wrote that... whoopsies.

OpenStudy (anonymous):

But then what?

OpenStudy (johnweldon1993):

That would be your answer... What you had was to simplify \[\large 1 - sin^2(x)\] Well we just saw that \[\large 1 - sin^2(x) = cos^2(x)\] so your simplification will be \(\large cos^2(x)\)

OpenStudy (anonymous):

Oh... okay. I feel really unintelligent at the moment. THANKS!

OpenStudy (johnweldon1993):

Don't worry about it, you're very intelligent :) and anytime!